You are offered an investment that requires an up-front (time = 0) payment of £47,000. This...

  1. You are offered an investment that requires an up-front (time = 0) payment of £47,000. This Investment “A” will provide you with a one-time return of £65,000 in four years. What is the implicit annual return of Investment A?

    A.

    6.67%

    B.

    6.73%

    C.

    8.44%

    D.

    9.00%

    E.

    9.63%
Answer #1

Let rate of return be x%

At this rate;present value of return=47000

47000=65000/1.0x^4

1.0x^4=(65000/47000)

1.0x=(65000/47000)^(1/4)

x=1.0844-1

=8.44%(Approx).